1.95 atm, how many moles of argon are in the container s tiled with 25.0 mol of (a) What is the number of molecules of oxygen in the balloon? molecules (c) What is the rms speed of axygen in the bailoon? O Type here to He
Answer #1

using formulas

Average kinetic energy KE =(3/2)*k*T

From ideal gas equation , P*V=n*R*T

rms speed v = sqrt (3*R*T/M)

3) volume s the balloog (V)-ITY3 3 V- 4IT -130.ollom3 From ideal gas eguatiun em 乔ー(1-013xldm)(6.ollom) ท n= 0.4679孑ー0.468 mol 2603 xt (6.WJ Хі89